OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -2x + 9y = 56 9x – 9y = 0

OpenStudy (campbell_st):

add the 2 equations together to eliminate y -2x + 9y = 56 9x - 9y = 0 ------------- 7x = 56 x = 8 substitute into the 1st equation to find y -2(8) + 9y = 56 -16 + 9y = 56 9y = 72 y = 8 the solution is x = 8 and y = 8